OpenStudy (anonymous):

Express the domain of f(x)= square root of (4-3x-x^2)?

OpenStudy (anonymous):

\[f(x)=\sqrt{4-3x-x^2}\] \[=\sqrt{(4+x )(1-x)}\] What can't x be? You know that the number inside the square root cannot be negative.

OpenStudy (anonymous):

What is the smallest x-value that DOES NOT make the number inside the square root sign a negative?

OpenStudy (anonymous):

You have the number 0 inside the square root sign but not less than that or else it will be a negative.

OpenStudy (anonymous):

I need to express the domain of f(x) =\[\sqrt{4-3x-x^2}\] in interval notation? It's for calculus

OpenStudy (anonymous):

We're taking it step by step. You're after this right? \[x>?\]

OpenStudy (anonymous):

We're taking it step by step here mate.

OpenStudy (anonymous):

What x value makes the whole term 0?

OpenStudy (anonymous):

@Jilian

OpenStudy (anonymous):

1 makes the whole term zero

OpenStudy (anonymous):

Good so x has to be greater than or equal to 1.

OpenStudy (anonymous):

\[x \ge1\]

OpenStudy (anonymous):

Do you understand these concepts of intervals now?

OpenStudy (anonymous):

If you go below 1 you will find yourself having a value less than 0 inside the square root which you don't want.

OpenStudy (anonymous):

Anything greater than one will make the solution non existing. I'm not clear on this interval notation however.

OpenStudy (anonymous):

less than one sorry

OpenStudy (anonymous):

My bad. I thought it was x-1 rather than 1-x

OpenStudy (anonymous):

Apologies.

OpenStudy (anonymous):

\[x \le1\]

OpenStudy (anonymous):

Now for the 4+x

OpenStudy (anonymous):

What's the maximum negative number you can put in their that won't make it negative. SO what number in there would make the value become zero?

OpenStudy (anonymous):

-4 makes ( 4+x) = 0

OpenStudy (anonymous):

Good work. so then x can't be less than -4. So then, x has to be greater than or equal to -4. \[x\ge -4\]

OpenStudy (anonymous):

Now you have these two inequalties. \[x\le 1\] \[x\ge -4\] Is x ranging between these two inequalities?

OpenStudy (anonymous):

so the domain will look like this, \[-4\le x \le1\]

OpenStudy (anonymous):

You owned it. Well Done. Excellent work mate.

OpenStudy (anonymous):

Thank You very much for your help. It was a pleasure working with you. You should be a Math professor.

OpenStudy (anonymous):

Now worries mate, and hahah, nah, I don't think so.

OpenStudy (anonymous):

No worries*